If 28.70 mL of 0.0200 M KMnO_4 is required to titr
Answer #1

The molar mass of complex is 491g/mol

with permanganate ion the oxalate is oxidised to CO2

thus 3C2O4-2 ions oxidise to 6CO2 and hence 6 electrons are given (6x3 ---> 6x4)

when permanganate is reduced to Mn+2 5 electrons are added.

The equivalents of KMNO4 = equivalents of complex. Let us assume the amount of oxalate in the complex is xg

28.7x10-3 x0.02x5 = xg /(491/6)

then x = 0.234g

Thus percentage of oxalate in the sample = (0.234/0.250)100 =93.6%